how many times should a restriction enzyme with a recognition of TCGA cut a DNA sequence of 1000 bases?

a) 2

b)4

c) 8

d) 16

e) 32
Answer #1

The answer is b) 4.

The steps inolved here is,

Step:1: Probability of occurnce of restriction site for TCGA sequence:

As one nitrogen base can occupay in one nucleotide position out of four, the probability of each base is 1/4.

So, the probability of TCGA sequence is 1/4*1/4*1/4*1/4= 1/256.

Step 2: Finding the number of restriction sites for 1000 bases DNA sequence=

No of bases in DNA sequences * probability of restriction site for TCGA sequence

= 1000*1/256 = 3.90 = ~4.

So, the answer is 4.